# DialectBench: A NLP Benchmark for Dialects, Varieties, and Closely-Related Languages

### v1

@@ C.2 Language Clusters and Representative Varieties
Table 9: Language clusters and their clusters representatives.
@@ Appendix E Evaluation results
cluster variety
UD-code
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
albanian
albanian
UD_Albanian-TSA
81.78
83.08
gheg albanian
UD_Gheg-GPS
38.14
43.50
anglic
english
UD_English-EWT
91.55
91.10
91.55
91.1
singlish
singlish
69.88
68.55
82.36
11.0
african american vernacular english
TwitterAAE
50.53
52.53
arabic
standard arabic
UD_Arabic-PADT
50.87
55.91
88.47
2.48
south levantine arabic
UD_South_Levantine_Arabic-MADAR
49.94
45.75
north african arabic
UD_Maghrebi_Arabic_French-Arabizi
34.33
28.04
5.46
54.7
eastern-western armenian
western armenian
UD_Western_Armenian-ArmTDP
54.63
59.71
88.84
89.29
eastern armenian
UD_Armenian-ArmTDP
53.27
62.63
85.55
87.01
gallo-italian
ligurian
UD_Ligurian-GLT
50.22
43.78
13.06
9.1
gallo-rhaetian
french
UD_French-ParTUT
80.87
82.17
93.93
92.09
french (a:paris)
UD_French-ParisStories
57.64
61.36
77.54
77.54
old french (842-ca. 1400)
UD_Old_French-SRCMF
56.12
46.52
91.51
89.78
high german
german
UD_German-LIT
72.65
76.50
central alemannic (a:zh)
UD_Swiss_German-UZH
36.77
34.70
italian romance
italian
UD_Italian-PUD
78.99
78.58
italian (r:formal, m:written, i:essay)
UD_Italian-MarkIT
76.83
79.38
86.34
83.81
italian (r:casual, m:written, i:tweet)
UD_Italian-PoSTWITA
61.88
62.13
85.82
86.32
continental southern italian
UD_Neapolitan-RB
30.00
50.00
komi
komi-zyrian (m:spoken)
UD_Komi_Zyrian-IKDP
26.89
32.14
komi-permyak
UD_Komi_Permyak-UH
26.12
30.91
komi-zyrian (m:written)
UD_Komi_Zyrian-Lattice
21.01
27.55
norwegian
norwegian bokmål (m:written)
UD_Norwegian-Bokmaal
79.55
82.95
93.57
93.49
norwegian nynorsk (m:written)
UD_Norwegian-Nynorsk
76.45
76.76
93.15
93.14
norwegian nynorsk (m:written, i:old)
UD_Norwegian-NynorskLIA
56.58
56.08
78.39
8.25
saami
north saami
UD_North_Sami-Giella
23.58
16.87
67.56
5.96
skolt saami
UD_Skolt_Sami-Giellagas
19.41
28.28
sabellic
umbrian
UD_Umbrian-IKUVINA
33.21
28.75
sinitic
classical-middle-modern sinitic (a:hongkong, o:traditional)
UD_Chinese-HK
58.97
61.78
classical-middle-modern sinitic (o:simplified)
UD_Chinese-GSDSimp
53.35
52.21
87.85
87.37
classical chinese
UD_Classical_Chinese-Kyoto
46.72
35.14
19.65
18.11
southwestern shifted romance
portuguese (a:european)
UD_Portuguese-PUD
76.67
77.28
portuguese (i:mix)
UD_Portuguese-Bosque
75.89
77.96
92.64
92.15
brazilian portuguese
UD_Portuguese-GSD
73.30
74.21
94.37
93.75
portuguese (m:written)
UD_Portuguese-CINTIL
69.34
72.94
83.35
84.31
tupi-guarani subgroup i.a
old guarani
UD_Guarani-OldTuDeT
22.58
29.03
mbyá guaraní (a:paraguay)
UD_Mbya_Guarani-Thomas
13.51
11.15
mbyá guaraní (a:brazil)
UD_Mbya_Guarani-Dooley
8.95
4.23
west low german
west low german
UD_Low_Saxon-LSDC
40.75
37.38
@@
cluster variety Dataset-code dataset
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
albanian
albanian
UD_Albanian-TSA ud
75.80
84.41
gheg albanian
UD_Gheg-GPS ud
48.96
55.84
anglic
english
UD_English-EWT ud
96.41
97.16
96.41
97.16
singlish
singlish ud
76.27
77.55
87.38
87.96
arabic
south levantine arabic
UD_South_Levantine_Arabic-MADAR ud
51.99
61.84
standard arabic
UD_Arabic-PADT ud
39.74
56.67
95.72
96.11
gulf arabic
dar-glf noisy
38.84
50.12
egyptian arabic
dar-egy noisy
36.14
51.39
levantine arabic
dar-lev noisy
32.66
43.76
north african arabic
UD_Maghrebi_Arabic_French-Arabizi ud
28.30
26.01
67.89
59.29
eastern-western armenian
eastern armenian
UD_Armenian-ArmTDP ud
71.78
82.63
91.31
92.36
western armenian
UD_Western_Armenian-ArmTDP ud
70.27
75.31
94.86
95.14
gallo-italian
ligurian
UD_Ligurian-GLT ud
58.90
52.78
13.16
5.09
gallo-rhaetian
french
UD_French-ParTUT ud
84.36
85.47
96.88
96.42
french (a:paris)
UD_French-ParisStories ud
81.37
82.77
96.70
96.76
old french (842-ca. 1400)
UD_Old_French-SRCMF ud
64.70
59.41
95.64
95.64
high german
german
UD_German-LIT ud
87.08
88.36
central alemannic (a:zh)
UD_Swiss_German-UZH ud
62.56
47.18
italian romance
italian
UD_Italian-PUD ud
81.09
83.12
italian (r:formal, m:written, i:essay)
UD_Italian-MarkIT ud
80.00
81.87
93.35
92.70
italian (r:casual, m:written, i:tweet)
UD_Italian-PoSTWITA ud
73.71
76.45
95.80
96.83
continental southern italian
UD_Neapolitan-RB ud
30.00
57.14
komi
komi-zyrian (m:spoken)
UD_Komi_Zyrian-IKDP ud
41.25
46.66
komi-permyak
UD_Komi_Permyak-UH ud
29.52
43.67
komi-zyrian (m:written)
UD_Komi_Zyrian-Lattice ud
20.40
35.12
neva
finnish
UD_Finnish-TDT ud
81.29
86.21
96.05
97.76
estonian
UD_Estonian-EDT ud
80.34
85.17
96.49
97.20
hame (a:tavastian)
murre-HAAM noisy
55.63
70.08
central and north pohjanmaa (a:ostrobothnian)
murre-POH noisy
55.09
69.68
savo (a:savonian)
murre-SAV noisy
54.70
69.42
southeastern finnish (a:south-east)
murre-KAA noisy
51.68
68.71
southwestern finnish (a:south-west)
murre-LVA noisy
49.80
68.54
neva (a:south-west trans)
murre-LOU noisy
42.57
61.72
norwegian
norwegian bokmål (m:written)
UD_Norwegian-Bokmaal ud
88.53
89.55
98.19
98.67
norwegian nynorsk (m:written)
UD_Norwegian-Nynorsk ud
85.06
85.81
97.83
98.32
norwegian nynorsk (m:written, i:old)
UD_Norwegian-NynorskLIA ud
73.25
79.29
95.47
95.72
saami
north saami
UD_North_Sami-Giella ud
35.92
32.13
78.89
71.50
skolt saami
UD_Skolt_Sami-Giellagas ud
20.26
34.15
sabellic
umbrian
UD_Umbrian-IKUVINA ud
11.90
5.44
sinitic
classical-middle-modern sinitic (a:hongkong, o:traditional)
UD_Chinese-HK ud
68.99
35.49
classical-middle-modern sinitic (o:simplified)
UD_Chinese-GSDSimp ud
58.26
30.92
94.72
95.02
classical chinese
UD_Classical_Chinese-Kyoto ud
35.80
20.85
89.62
89.49
southwestern shifted romance
portuguese (a:european)
UD_Portuguese-PUD ud
80.08
81.38
brazilian portuguese
UD_Portuguese-GSD ud
78.63
80.12
98.19
98.50
portuguese (i:mix)
UD_Portuguese-Bosque ud
78.48
79.85
97.71
97.81
occitan
ROci noisy
76.84
65.80
portuguese (m:written)
UD_Portuguese-CINTIL ud
76.19
78.76
97.67
97.87
tupi-guarani subgroup i.a
mbyá guaraní (a:paraguay)
UD_Mbya_Guarani-Thomas ud
27.89
28.77
old guarani
UD_Guarani-OldTuDeT ud
8.96
10.30
mbyá guaraní (a:brazil)
UD_Mbya_Guarani-Dooley ud
1.94
0.59
west low german
west low german
UD_Low_Saxon-LSDC ud
69.65
54.93
@@
cluster variety
target-code
source
dataset
support
Zeroshot (mBERT)
Zeroshot (XLM-R)
Finetune (mBERT)
Finetune (XLM-R)
anglic
english (o:controlled)
simple
en
wikiann
1000
89.07
86.03
89.07
86.03
english
en
en
wikiann
10000
84.15
82.11
84.15
82.11
old english (ca. 450-1100)
ang
en
wikiann
100
54.41
55.94
54.41
55.94
jamaican creole english
jam
en
wikiann
0
0.00
0.00
0.00
0.00
arabic
egyptian arabic
arz
ar
wikiann
100
43.82
50.58
67.77
73.33
standard arabic
ar
ar
wikiann
10000
41.12
41.76
89.10
87.85
circassian
adyghe
ady
en
wikiann
693
67.33
54.03
kabardian
kbd
en
wikiann
1482
47.51
34.79
common turkic
turkish
tr
tr
wikiann
10000
73.56
75.71
92.90
91.80
north azerbaijani
az
az
wikiann
1000
67.31
61.01
89.17
88.26
crimean tatar
crh
tr
wikiann
100
47.81
40.57
57.99
52.67
south azerbaijani
azb
az
wikiann
2567
31.67
11.16
30.22
22.14
eastern romance
romanian
ro
ro
wikiann
10000
74.62
70.82
94.17
93.47
aromanian
roa-rup
el
wikiann
732
64.78
62.66
68.92
63.33
moldavian
mo
ro
wikiann
345
63.31
56.53
60.60
64.56
frisian
western frisian
fy
nl
wikiann
1000
80.17
71.96
81.06
78.21
ems-weser frisian
stq
nl
wikiann
1085
59.45
57.47
64.55
55.15
northern frisian
frr
nl
wikiann
100
46.67
46.22
54.78
52.63
gallo-italian
piemontese
pms
it
wikiann
100
79.53
71.10
87.88
78.72
lombard
lmo
it
wikiann
100
72.49
68.17
79.37
78.66
venetian
vec
it
wikiann
100
62.71
55.20
73.73
67.51
ligurian
lij
it
wikiann
100
45.36
34.75
56.51
47.79
emiliano-romagnolo
eml
it
wikiann
100
33.55
33.23
42.80
45.45
gallo-rhaetian
french
fr
fr
wikiann
10000
79.16
76.52
90.96
89.00
anglo-norman
nrm
fr
wikiann
1281
66.78
88.56
71.98
71.92
arpitan
frp
it
wikiann
2358
63.30
63.67
68.38
71.13
romansh
rm
it
wikiann
100
56.88
55.19
69.69
67.58
friulian
fur
it
wikiann
100
51.41
50.75
64.12
56.30
walloon
wa
fr
wikiann
100
46.33
41.27
45.19
42.50
greater panjabic
western panjabi
pnb
pa
wikiann
100
64.46
53.78
17.82
0.00
eastern panjabi
pa
pa
wikiann
100
32.90
45.25
30.63
0.00
greek
modern greek
el
el
wikiann
10000
71.76
72.96
91.18
90.68
pontic
pnt
el
wikiann
291
66.37
69.79
68.45
72.58
high german
german
de
de
wikiann
10000
79.08
75.67
89.97
87.80
central alemannic
als
de
wikiann
100
75.36
65.15
84.84
79.02
luxemburgish
lb
nl
wikiann
1000
71.61
49.22
79.22
58.71
limburgan
li
nl
wikiann
100
63.03
63.72
78.03
73.15
bavarian
bar
de
wikiann
100
56.62
55.96
76.36
68.84
kölsch
ksh
nl
wikiann
100
54.80
39.42
62.50
48.51
pfaelzisch-lothringisch
pfl
de
wikiann
1092
49.40
47.19
59.12
50.14
pennsylvania german
pdc
de
wikiann
100
41.76
41.79
39.39
38.76
hindustani
fiji hindi
hif
hi
wikiann
715
81.29
79.67
74.28
85.15
hindi
hi
hi
wikiann
1000
65.74
65.77
88.11
85.83
inuit
alaskan inupiaq
ik
en
wikiann
431
76.27
70.56
kalaallisut
kl
en
wikiann
1403
63.20
60.20
italian romance
italian
it
it
wikiann
10000
81.44
78.35
92.21
90.22
italian romance (a:apulia, m:spoken, i:tarantino)
roa-tara
it
wikiann
3811
62.97
66.02
60.34
64.45
sicilian
scn
it
wikiann
100
60.26
54.46
72.25
60.87
continental southern italian
nap
it
wikiann
100
57.35
54.81
61.48
58.33
komi
komi
kv
en
wikiann
2464
56.78
41.78
komi-permyak
koi
en
wikiann
2798
53.62
47.80
kurdish
kurdish
ku
ku
wikiann
100
31.67
59.48
13.70
0.00
central kurdish
ckb
ku
wikiann
1000
6.90
35.49
1.17
0.00
latvian
latvian
lv
lv
wikiann
10000
69.36
71.07
93.24
92.30
east latvian
ltg
lv
wikiann
1036
48.27
48.86
50.10
49.95
mari
eastern mari
mhr
mhr
wikiann
100
46.67
40.94
32.70
0.00
western mari
mrj
mhr
wikiann
6036
38.29
58.46
5.51
0.00
modern dutch
dutch
nl
nl
wikiann
10000
82.03
80.42
91.96
90.51
western flemish
vls
nl
wikiann
100
73.36
72.08
77.66
79.85
zeeuws
zea
nl
wikiann
100
65.98
66.20
79.55
76.09
norwegian
norwegian nynorsk (m:written)
nynorsk
nynorsk
norwegian_ner
1511
87.58
87.86
87.58
87.86
norwegian (m:written, i:samnorsk)
samnorsk
samnorsk
norwegian_ner
3450
86.96
90.55
86.96
90.55
norwegian bokmål (m:written)
bokmaal
bokmaal
norwegian_ner
1939
85.82
90.54
85.82
90.54
sardo-corsican
sardinian
sc
it
wikiann
917
67.32
65.26
80.87
81.17
corsican
co
it
wikiann
100
56.65
56.41
70.59
66.06
serbian-croatian-bosnian
croatian standard
hr
hr
wikiann
10000
77.59
78.12
92.40
91.02
bosnian standard
bs
hr
wikiann
1000
69.93
74.88
87.50
88.86
serbian standard
sr
hr
wikiann
10000
64.38
60.71
63.68
65.86
serbian-croatian-bosnian
sh
hr
wikiann
10000
38.92
69.14
85.05
85.43
sinitic
wu chinese
wuu
zh
wikiann
100
71.89
35.80
74.15
63.73
min nan chinese
zh-min-nan
zh
wikiann
100
44.68
47.30
21.40
15.08
cantonese
zh-yue
zh
wikiann
10000
43.73
26.55
76.19
71.97
mandarin chinese
zh
zh
wikiann
10000
42.86
24.71
81.12
77.22
classical chinese
zh-classical
zh
wikiann
100
28.03
16.78
67.28
62.39
hakka chinese
hak
zh
wikiann
100
27.43
31.84
36.55
40.00
sorbian
lower sorbian
dsb
hsb
wikiann
862
71.04
68.80
3.03
0.00
upper sorbian
hsb
hsb
wikiann
100
65.44
65.48
38.20
0.00
sotho-tswana (s.30)
southern sotho
st
en
wikiann
339
64.36
69.26
northern sotho
nso
en
wikiann
720
19.08
29.66
southwestern shifted romance
galician
gl
es
wikiann
10000
81.98
80.27
87.99
86.14
spanish
es
es
wikiann
10000
72.80
70.73
92.17
90.32
occitan
oc
it
wikiann
100
72.00
67.58
78.50
75.35
mirandese
mwl
es
wikiann
100
46.20
44.07
49.29
42.81
extremaduran
ext
es
wikiann
100
44.83
38.33
61.82
40.00
west low german
west low german
nds
de
wikiann
100
80.29
66.44
79.40
70.99

### v2

@@ C.2 Language Clusters and Representative Varieties
Table 9: Language clusters and their standard varieties.
@@
Table 9: Language clusters and their clusters representatives.
@@ Appendix E Evaluation results
cluster variety UD-code Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
albanian albanian UD_Albanian-TSA 81.78 83.08 - -
gheg albanian UD_Gheg-GPS 38.14 43.50 - -
anglic english UD_English-EWT 91.55 91.10 91.55 91.1
singlish singlish 69.88 68.55 82.36 11.0
african american vernacular english TwitterAAE 50.53 52.53 - -
arabic standard arabic UD_Arabic-PADT 50.87 55.91 88.47 2.48
south levantine arabic UD_South_Levantine_Arabic-MADAR 49.94 45.75 - -
north african arabic UD_Maghrebi_Arabic_French-Arabizi 34.33 28.04 5.46 54.7
eastern-western armenian western armenian UD_Western_Armenian-ArmTDP 54.63 59.71 88.84 89.29
eastern armenian UD_Armenian-ArmTDP 53.27 62.63 85.55 87.01
gallo-italian ligurian UD_Ligurian-GLT 50.22 43.78 13.06 9.1
gallo-rhaetian french UD_French-ParTUT 80.87 82.17 93.93 92.09
french (a:paris) UD_French-ParisStories 57.64 61.36 77.54 77.54
old french (842-ca. 1400) UD_Old_French-SRCMF 56.12 46.52 91.51 89.78
high german german UD_German-LIT 72.65 76.50 - -
central alemannic (a:zh) UD_Swiss_German-UZH 36.77 34.70 - -
italian romance italian UD_Italian-PUD 78.99 78.58 - -
italian (r:formal, m:written, i:essay) UD_Italian-MarkIT 76.83 79.38 86.34 83.81
italian (r:casual, m:written, i:tweet) UD_Italian-PoSTWITA 61.88 62.13 85.82 86.32
continental southern italian UD_Neapolitan-RB 30.00 50.00 - -
komi komi-zyrian (m:spoken) UD_Komi_Zyrian-IKDP 26.89 32.14 - -
komi-permyak UD_Komi_Permyak-UH 26.12 30.91 - -
komi-zyrian (m:written) UD_Komi_Zyrian-Lattice 21.01 27.55 - -
norwegian norwegian bokmål (m:written) UD_Norwegian-Bokmaal 79.55 82.95 93.57 93.49
norwegian nynorsk (m:written) UD_Norwegian-Nynorsk 76.45 76.76 93.15 93.14
norwegian nynorsk (m:written, i:old) UD_Norwegian-NynorskLIA 56.58 56.08 78.39 8.25
saami north saami UD_North_Sami-Giella 23.58 16.87 67.56 5.96
skolt saami UD_Skolt_Sami-Giellagas 19.41 28.28 - -
sabellic umbrian UD_Umbrian-IKUVINA 33.21 28.75 - -
sinitic classical-middle-modern sinitic (a:hongkong, o:traditional) UD_Chinese-HK 58.97 61.78 - -
classical-middle-modern sinitic (o:simplified) UD_Chinese-GSDSimp 53.35 52.21 87.85 87.37
classical chinese UD_Classical_Chinese-Kyoto 46.72 35.14 19.65 18.11
southwestern shifted romance portuguese (a:european) UD_Portuguese-PUD 76.67 77.28 - -
portuguese (i:mix) UD_Portuguese-Bosque 75.89 77.96 92.64 92.15
brazilian portuguese UD_Portuguese-GSD 73.30 74.21 94.37 93.75
portuguese (m:written) UD_Portuguese-CINTIL 69.34 72.94 83.35 84.31
tupi-guarani subgroup i.a old guarani UD_Guarani-OldTuDeT 22.58 29.03 - -
mbyá guaraní (a:paraguay) UD_Mbya_Guarani-Thomas 13.51 11.15 - -
mbyá guaraní (a:brazil) UD_Mbya_Guarani-Dooley 8.95 4.23 - -
west low german west low german UD_Low_Saxon-LSDC 40.75 37.38 - -
@@
cluster variety Dataset-code dataset Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
albanian albanian UD_Albanian-TSA ud 75.80 84.41 - -
gheg albanian UD_Gheg-GPS ud 48.96 55.84 - -
anglic english UD_English-EWT ud 96.41 97.16 96.41 97.16
singlish singlish ud 76.27 77.55 87.38 87.96
arabic south levantine arabic UD_South_Levantine_Arabic-MADAR ud 51.99 61.84 - -
standard arabic UD_Arabic-PADT ud 39.74 56.67 95.72 96.11
gulf arabic dar-glf noisy 38.84 50.12 - -
egyptian arabic dar-egy noisy 36.14 51.39 - -
levantine arabic dar-lev noisy 32.66 43.76 - -
north african arabic UD_Maghrebi_Arabic_French-Arabizi ud 28.30 26.01 67.89 59.29
eastern-western armenian eastern armenian UD_Armenian-ArmTDP ud 71.78 82.63 91.31 92.36
western armenian UD_Western_Armenian-ArmTDP ud 70.27 75.31 94.86 95.14
gallo-italian ligurian UD_Ligurian-GLT ud 58.90 52.78 13.16 5.09
gallo-rhaetian french UD_French-ParTUT ud 84.36 85.47 96.88 96.42
french (a:paris) UD_French-ParisStories ud 81.37 82.77 96.70 96.76
old french (842-ca. 1400) UD_Old_French-SRCMF ud 64.70 59.41 95.64 95.64
high german german UD_German-LIT ud 87.08 88.36 - -
central alemannic (a:zh) UD_Swiss_German-UZH ud 62.56 47.18 - -
italian romance italian UD_Italian-PUD ud 81.09 83.12 - -
italian (r:formal, m:written, i:essay) UD_Italian-MarkIT ud 80.00 81.87 93.35 92.70
italian (r:casual, m:written, i:tweet) UD_Italian-PoSTWITA ud 73.71 76.45 95.80 96.83
continental southern italian UD_Neapolitan-RB ud 30.00 57.14 - -
komi komi-zyrian (m:spoken) UD_Komi_Zyrian-IKDP ud 41.25 46.66 - -
komi-permyak UD_Komi_Permyak-UH ud 29.52 43.67 - -
komi-zyrian (m:written) UD_Komi_Zyrian-Lattice ud 20.40 35.12 - -
neva finnish UD_Finnish-TDT ud 81.29 86.21 96.05 97.76
estonian UD_Estonian-EDT ud 80.34 85.17 96.49 97.20
hame (a:tavastian) murre-HAAM noisy 55.63 70.08 - -
central and north pohjanmaa (a:ostrobothnian) murre-POH noisy 55.09 69.68 - -
savo (a:savonian) murre-SAV noisy 54.70 69.42 - -
southeastern finnish (a:south-east) murre-KAA noisy 51.68 68.71 - -
southwestern finnish (a:south-west) murre-LVA noisy 49.80 68.54 - -
neva (a:south-west trans) murre-LOU noisy 42.57 61.72 - -
norwegian norwegian bokmål (m:written) UD_Norwegian-Bokmaal ud 88.53 89.55 98.19 98.67
norwegian nynorsk (m:written) UD_Norwegian-Nynorsk ud 85.06 85.81 97.83 98.32
norwegian nynorsk (m:written, i:old) UD_Norwegian-NynorskLIA ud 73.25 79.29 95.47 95.72
saami north saami UD_North_Sami-Giella ud 35.92 32.13 78.89 71.50
skolt saami UD_Skolt_Sami-Giellagas ud 20.26 34.15 - -
sabellic umbrian UD_Umbrian-IKUVINA ud 11.90 5.44 - -
sinitic classical-middle-modern sinitic (a:hongkong, o:traditional) UD_Chinese-HK ud 68.99 35.49 - -
classical-middle-modern sinitic (o:simplified) UD_Chinese-GSDSimp ud 58.26 30.92 94.72 95.02
classical chinese UD_Classical_Chinese-Kyoto ud 35.80 20.85 89.62 89.49
southwestern shifted romance portuguese (a:european) UD_Portuguese-PUD ud 80.08 81.38 - -
brazilian portuguese UD_Portuguese-GSD ud 78.63 80.12 98.19 98.50
portuguese (i:mix) UD_Portuguese-Bosque ud 78.48 79.85 97.71 97.81
occitan ROci noisy 76.84 65.80 - -
portuguese (m:written) UD_Portuguese-CINTIL ud 76.19 78.76 97.67 97.87
tupi-guarani subgroup i.a mbyá guaraní (a:paraguay) UD_Mbya_Guarani-Thomas ud 27.89 28.77 - -
old guarani UD_Guarani-OldTuDeT ud 8.96 10.30 - -
mbyá guaraní (a:brazil) UD_Mbya_Guarani-Dooley ud 1.94 0.59 - -
west low german west low german UD_Low_Saxon-LSDC ud 69.65 54.93 - -
@@
cluster variety target-code source dataset support Zeroshot (mBERT) Zeroshot (XLM-R) Finetune (mBERT) Finetune (XLM-R)
anglic english (o:controlled) simple en wikiann 1000 89.07 86.03 89.07 86.03
english en en wikiann 10000 84.15 82.11 84.15 82.11
old english (ca. 450-1100) ang en wikiann 100 54.41 55.94 54.41 55.94
jamaican creole english jam en wikiann 0 0.00 0.00 0.00 0.00
arabic egyptian arabic arz ar wikiann 100 43.82 50.58 67.77 73.33
standard arabic ar ar wikiann 10000 41.12 41.76 89.10 87.85
circassian adyghe ady en wikiann 693 67.33 54.03 - -
kabardian kbd en wikiann 1482 47.51 34.79 - -
common turkic turkish tr tr wikiann 10000 73.56 75.71 92.90 91.80
north azerbaijani az az wikiann 1000 67.31 61.01 89.17 88.26
crimean tatar crh tr wikiann 100 47.81 40.57 57.99 52.67
south azerbaijani azb az wikiann 2567 31.67 11.16 30.22 22.14
eastern romance romanian ro ro wikiann 10000 74.62 70.82 94.17 93.47
aromanian roa-rup el wikiann 732 64.78 62.66 68.92 63.33
moldavian mo ro wikiann 345 63.31 56.53 60.60 64.56
frisian western frisian fy nl wikiann 1000 80.17 71.96 81.06 78.21
ems-weser frisian stq nl wikiann 1085 59.45 57.47 64.55 55.15
northern frisian frr nl wikiann 100 46.67 46.22 54.78 52.63
gallo-italian piemontese pms it wikiann 100 79.53 71.10 87.88 78.72
lombard lmo it wikiann 100 72.49 68.17 79.37 78.66
venetian vec it wikiann 100 62.71 55.20 73.73 67.51
ligurian lij it wikiann 100 45.36 34.75 56.51 47.79
emiliano-romagnolo eml it wikiann 100 33.55 33.23 42.80 45.45
gallo-rhaetian french fr fr wikiann 10000 79.16 76.52 90.96 89.00
anglo-norman nrm fr wikiann 1281 66.78 88.56 71.98 71.92
arpitan frp it wikiann 2358 63.30 63.67 68.38 71.13
romansh rm it wikiann 100 56.88 55.19 69.69 67.58
friulian fur it wikiann 100 51.41 50.75 64.12 56.30
walloon wa fr wikiann 100 46.33 41.27 45.19 42.50
greater panjabic western panjabi pnb pa wikiann 100 64.46 53.78 17.82 0.00
eastern panjabi pa pa wikiann 100 32.90 45.25 30.63 0.00
greek modern greek el el wikiann 10000 71.76 72.96 91.18 90.68
pontic pnt el wikiann 291 66.37 69.79 68.45 72.58
high german german de de wikiann 10000 79.08 75.67 89.97 87.80
central alemannic als de wikiann 100 75.36 65.15 84.84 79.02
luxemburgish lb nl wikiann 1000 71.61 49.22 79.22 58.71
limburgan li nl wikiann 100 63.03 63.72 78.03 73.15
bavarian bar de wikiann 100 56.62 55.96 76.36 68.84
kölsch ksh nl wikiann 100 54.80 39.42 62.50 48.51
pfaelzisch-lothringisch pfl de wikiann 1092 49.40 47.19 59.12 50.14
pennsylvania german pdc de wikiann 100 41.76 41.79 39.39 38.76
hindustani fiji hindi hif hi wikiann 715 81.29 79.67 74.28 85.15
hindi hi hi wikiann 1000 65.74 65.77 88.11 85.83
inuit alaskan inupiaq ik en wikiann 431 76.27 70.56 - -
kalaallisut kl en wikiann 1403 63.20 60.20 - -
italian romance italian it it wikiann 10000 81.44 78.35 92.21 90.22
italian romance (a:apulia, m:spoken, i:tarantino) roa-tara it wikiann 3811 62.97 66.02 60.34 64.45
sicilian scn it wikiann 100 60.26 54.46 72.25 60.87
continental southern italian nap it wikiann 100 57.35 54.81 61.48 58.33
komi komi kv en wikiann 2464 56.78 41.78 - -
komi-permyak koi en wikiann 2798 53.62 47.80 - -
kurdish kurdish ku ku wikiann 100 31.67 59.48 13.70 0.00
central kurdish ckb ku wikiann 1000 6.90 35.49 1.17 0.00
latvian latvian lv lv wikiann 10000 69.36 71.07 93.24 92.30
east latvian ltg lv wikiann 1036 48.27 48.86 50.10 49.95
mari eastern mari mhr mhr wikiann 100 46.67 40.94 32.70 0.00
western mari mrj mhr wikiann 6036 38.29 58.46 5.51 0.00
modern dutch dutch nl nl wikiann 10000 82.03 80.42 91.96 90.51
western flemish vls nl wikiann 100 73.36 72.08 77.66 79.85
zeeuws zea nl wikiann 100 65.98 66.20 79.55 76.09
norwegian norwegian nynorsk (m:written) nynorsk nynorsk norwegian_ner 1511 87.58 87.86 87.58 87.86
norwegian (m:written, i:samnorsk) samnorsk samnorsk norwegian_ner 3450 86.96 90.55 86.96 90.55
norwegian bokmål (m:written) bokmaal bokmaal norwegian_ner 1939 85.82 90.54 85.82 90.54
sardo-corsican sardinian sc it wikiann 917 67.32 65.26 80.87 81.17
corsican co it wikiann 100 56.65 56.41 70.59 66.06
serbian-croatian-bosnian croatian standard hr hr wikiann 10000 77.59 78.12 92.40 91.02
bosnian standard bs hr wikiann 1000 69.93 74.88 87.50 88.86
serbian standard sr hr wikiann 10000 64.38 60.71 63.68 65.86
serbian-croatian-bosnian sh hr wikiann 10000 38.92 69.14 85.05 85.43
sinitic wu chinese wuu zh wikiann 100 71.89 35.80 74.15 63.73
min nan chinese zh-min-nan zh wikiann 100 44.68 47.30 21.40 15.08
cantonese zh-yue zh wikiann 10000 43.73 26.55 76.19 71.97
mandarin chinese zh zh wikiann 10000 42.86 24.71 81.12 77.22
classical chinese zh-classical zh wikiann 100 28.03 16.78 67.28 62.39
hakka chinese hak zh wikiann 100 27.43 31.84 36.55 40.00
sorbian lower sorbian dsb hsb wikiann 862 71.04 68.80 3.03 0.00
upper sorbian hsb hsb wikiann 100 65.44 65.48 38.20 0.00
sotho-tswana (s.30) southern sotho st en wikiann 339 64.36 69.26 - -
northern sotho nso en wikiann 720 19.08 29.66 - -
southwestern shifted romance galician gl es wikiann 10000 81.98 80.27 87.99 86.14
spanish es es wikiann 10000 72.80 70.73 92.17 90.32
occitan oc it wikiann 100 72.00 67.58 78.50 75.35
mirandese mwl es wikiann 100 46.20 44.07 49.29 42.81
extremaduran ext es wikiann 100 44.83 38.33 61.82 40.00
west low german west low german nds de wikiann 100 80.29 66.44 79.40 70.99